# Tip growth in a strongly concentrated aggregation model follows local geodesics

### v2

@@ 3.1 ALE to auxiliary model
The two maps above whose derivatives we consider can be generated by the backward equation using driving functions which we will call $\xi^{0}$ (generating $\Phi^{\mathrm{ALE}}_{0}\circ\dots\circ\Psi_{N_{j}-1}^{j}$) and $\xi^{1}$ (generating $R_{\Delta_{1}^{n}}\circ\dots\circ\overline{\Psi}^{j}_{N_{j}-1}$). We can now use Lemma 2.5 to compare the above derivatives. We have the bound $\|\xi^{0}-\xi^{1}\|_{\infty}\leq|\delta_{1}|+\dots+|\delta_{n}|$. Taking $z_{0}=f^{\mathrm{ALE}}_{n_{j}(N_{j})}(\Psi_{N_{j}}(e^{\sigma+i\theta}e^{i\phi_{n}^{j}}))$, we have $|z_{0}|-1\geq\mathbf{c}^{1/2}$, so using standard estimates for conformal maps (see for example Section 1.1 of [13]), we have $\frac{\mathbf{c}^{1/2}}{A}\leq|(u_{t}^{0})^{\prime}(z_{0})|\leq\frac{A}{\mathbf{c}^{1/2}}$ for all $0<t<T$, where the constant $A$ depends only on $T$ and $K_{0}$. This gives us, modifying the constant $A=A(T,K_{0})$ appropriately (as we will throughout), $\frac{\mathbf{c}^{1/2}}{A}t\leq|\Lambda_{t}|\leq\frac{A}{\mathbf{c}^{3/2}}t$.
@@
$f(w)=1+d(\mathbf{c})+O\left(\frac{|w-1|^{2}}{\mathbf{c}^{1/2}}\right),$ (3.8)
@@
(see [14, equation (26)]) and using (3.5) and (3.8) we know
@@
and so by standard conformal map estimates,
@@
Thus
@@
We will classify the point $z=e^{\sigma+i\theta}$ based on its projection $\hat{z}=e^{i\theta}$ and follow them both through their backwards evolutions, so let $z_{k}=(f_{k}\circ f_{k+1}\circ\dots\circ f_{n})(z)$, and also let $\hat{z}_{k}=(f_{k}\circ f_{k+1}\circ\dots\circ f_{n})(\hat{z})$. Let $L$ be a lower bound on $\min_{l\leq n}\min_{i\not=j}|\phi_{l}^{i}-\phi_{l}^{j}|$, let $A_{L/2}$ be the constant in Lemma 2.4 corresponding to points further than $L/2$ from 1, and let $A_{T}=e^{A_{L/2}T}$. Observe that there can be at most two values of $k$ such that both $\hat{z}_{k}\in\mathbb{T}$ and $|\hat{z}_{k}-e^{i\theta^{\mathrm{ALE}}_{k-1}}|<|e^{i(\beta_{\mathbf{c}}+\frac{\mathbf{c}^{1/2}}{A_{T}})}-1|$ because after the second condition is met once, applying $\left\lfloor T/\mathbf{c}\right\rfloor$ maps cannot move $\hat{z}_{k}$ further than $\beta$ away from a tip, and it will be taken off the circle the next time a slit is attached near to it. This gives us three cases depending on the number of $k$s.
First suppose there is no such $k$. Then by Lemma 3.6, $|(f^{\mathrm{ALE}}_{1}\circ\dots\circ f^{\mathrm{ALE}}_{n})^{\prime}(\hat{z})|\leq B\mathbf{c}^{-1/2}$ almost surely, so by (2.7) from Lemma 2.5,
@@
This gives us $|(f^{\mathrm{ALE}}_{1}\circ\dots\circ f^{\mathrm{ALE}}_{n})(z)|-1\leq B\mathbf{c}^{-1/2}\sigma$, and hence by Lemma 3.5, $|(f^{\mathrm{ALE}}_{1}\circ\dots\circ f^{\mathrm{ALE}}_{n})^{\prime}(z)|>1.$ Let $j$ be the index of the closest $\phi_{0}^{j}$ to $(f^{\mathrm{ALE}}_{1}\circ\dots\circ f^{\mathrm{ALE}}_{n})(\hat{z})$. We can decompose $\Phi_{0}$ as $f_{t_{j},\phi_{0}^{j}}\circ\Psi^{j}_{0}$, where $\Psi^{j}_{0}$ is generated in the reverse equation by a driving function run for total time $c_{0}-t^{j}$ which stays at least distance $L/4$ from $(u_{s}(\hat{z}))_{s\leq c_{0}-t^{j}}$ at all times. Then using equation (26) of [14],
@@
Next, if there is only one such $k$, we must have $|\hat{z}_{k}-e^{i\theta^{\mathrm{ALE}}_{k-1}}|\geq\frac{\mathbf{c}^{1/2}}{2A_{T}^{2}},$ since there is an $l\geq k$ with $\theta^{\mathrm{ALE}}_{l}$ attached at the same slit as $\theta^{\mathrm{ALE}}_{k-1}$ and if $\hat{z}_{k}$ were any closer to $e^{i\theta^{\mathrm{ALE}}_{k-1}}$ then $\hat{z}_{l+1}$ would be within $|e^{i(\beta_{\mathbf{c}}+\frac{\mathbf{c}^{1/2}}{A_{T}})}-1|$ of $e^{i\theta^{\mathrm{ALE}}_{l}}$.
@@
and Lemma 2.3 tells us that $|z_{k-1}-e^{i\theta^{\mathrm{ALE}}_{k-1}}|\asymp\mathbf{c}^{1/4}\delta^{1/2}$. If the attachment point $\theta^{\mathrm{ALE}}_{k-1}$ is on the $j$th slit, then it is within $D$ of the preimage $\phi^{j}_{k-2}$ of $\theta^{\mathrm{ALE}}_{l}$ under $f^{\mathrm{ALE}}_{l+1}\circ\dots\circ f^{\mathrm{ALE}}_{k-2}$, where $l<k-1$ was the previous time a particle was attached at the $j$th slit (without loss of generality such an $l$ exists, as we can decompose the initial condition to have a slit of capacity $\mathbf{c}$ at the top of a longer slit in position $j$).
@@
If there are two values of $k$, $k_{2}<k_{1}$, we can find a bound of the same order using the same argument as above, replacing $k$ by $k_{1}$ and $l$ by $k_{2}$. ∎
@@ 3.2 Auxiliary to multinomial model
$\left|\log\frac{|(\Phi^{*}_{l})^{\prime}(e^{\sigma+i(\phi_{l}^{j}+\theta)})|}{|e^{\sigma+i\theta}-1|\times|(\Phi^{*}_{l})^{\prime\prime}(e^{i\phi_{l}^{j}})|}\right|$ $\leq\frac{|e^{\sigma+i\theta}-1|\times\sup_{z\in\gamma}|(\Phi^{*}_{l})^{(3)}(z)|}{|(\Phi^{*}_{l})^{\prime\prime}(e^{i\phi_{l}^{j}})|}.$ (3.18)
@@ 4 Multinomial to Laplacian path model
where =$\mathcal{H}=\{\varphi\in C(X):\|\varphi\|_{\mathrm{Lip}}+\|\varphi\|_{\infty}\leq 1\}$, for
@@
Under the same assumptions as Theorem 1.1, there exists a universal constant $R$ and a constant $A=A(T,L,k,K_{0},\eta)$ such that
@@
The latter two terms are small with high probability using Theorem 4.2, so we only need to bound the ”martingale” terms $\int_{0}^{T}(I^{j}_{t}-p^{j}_{t})\varphi(x_{j}(t))\,\mathrm{d}t$ for each $j$. We can write this as
@@
Since $|p^{j}_{n\mathbf{c}}-p^{j}_{t}|\leq A\mathbf{c}$ by Proposition 4.4, we have $|\sum_{n=0}^{\frac{T}{\mathbf{c}}-1}\int_{n\mathbf{c}}^{(n+1)\mathbf{c}}(p^{j}_{n\mathbf{c}}-p^{j}_{t})\varphi(x_{j}(t))\,\mathrm{d}t|\leq AT\mathbf{c}$. For the remaining term, we would like to simply compute second moments and so show the term is small, but need some way of doing this uniformly in $\varphi$. Choose some large $N$, and for simplicity assume it is a factor of $T/\mathbf{c}$. Then write
@@
Our proof of Theorem 4.2 has two parts: first we show in Proposition 4.4 that $t\mapsto(x_{j}(t),\bar{x}_{j}(t),p^{j}_{t},\bar{p}^{j}_{t})$ is Lipschitz on $[0,T]$, which will allow us to deduce global bounds on the $|x_{j}(t)-\bar{x}_{j}(t)|$ and $|p^{j}_{t}-\bar{p}^{j}_{t}|$ by showing that they are close at a sufficiently dense set of times.
@@ 4.1 Lipschitz bound on the locations and probabilities
There exists a constant $A=A(k,K_{0},T,\eta)$ such that, almost surely, for $0\leq t,s\leq T$, $|x_{j}(t)-x_{j}(s)|\leq A|t-s|$ and $|p_{t}^{j}-p_{s}^{j}|\leq A|t-s|$, for all $j=1,\dots,k$ in the multinomial model. For the Laplacian path model, for all $j$ and for $0\leq t,s\leq T$, we also have $|\bar{x}_{j}(t)-\bar{x}_{j}(s)|\leq A|t-s|$ and $|\bar{p}_{t}^{j}-\bar{p}_{s}^{j}|\leq A|t-s|$.
We prove this by carefully examining the derivatives of $\Phi_{t}$ and $\bar{\Phi}_{t}$ near the tips. The following lemma examines higher derivatives than necessary for this paper because it is not too much extra work and may be useful in future research.
Lemma 4.5.
Let $\Phi_{t}$ be the solution to Loewner’s equation for driving function $\xi_{t}$ corresponding to the ALE as above. Let $q_{t}(z)=\frac{\partial}{\partial t}\Phi_{t}(z)=\Phi_{t}^{\prime}(z)z\frac{z+\xi_{t}}{z-\xi_{t}}$ for $z\in\Delta$. Note that we can write Loewner’s equation in the integral form $\Phi_{t}(z)=\Phi_{0}(z)+\int_{0}^{t}q_{s}(z)\,\mathrm{d}s$ and for $m\geq 1$, $\Phi_{t}^{(m)}(z)=\Phi_{0}^{(m)}(z)+\int_{0}^{t}q_{s}^{(m)}(z)\,\mathrm{d}s$. Then for $m\geq 0$, at times $t$ when $\xi_{t}=x_{j}(t)$ for some $j$, we have
$q_{t}^{(m)}(x_{j}(t)):=\lim_{z\to x_{j}(t)}q_{t}^{(m)}(z)$ $=m\Phi_{t}^{(m)}(\xi_{t})+3\xi_{t}\Phi_{t}^{(m+1)}(\xi_{t})+\frac{2\xi_{t}^{2}}{m+1}\Phi_{t}^{(m+2)}(\xi_{t}).$
@@
Let $h(z)=z\frac{z+\xi_{t}}{z-\xi_{t}}=z+2\xi_{t}+\frac{2\xi_{t}^{2}}{z-\xi_{t}}$. Then $h^{\prime}(z)=1-\frac{2\xi_{t}^{2}}{(z-\xi_{t})^{2}}$, and for $p\geq 2$, $h^{(p)}(z)=2\xi_{t}^{2}\frac{(-1)^{p}p!}{(z-\xi_{t})^{p+1}}$.
Using the product rule,
$q_{t}^{(m)}(z)$ $=\sum_{p=0}^{m}\binom{m}{p}h^{(m-p)}(z)\Phi_{t}^{(p+1)}(z)$ (4.2)
$=2\xi_{t}^{2}\sum_{p=0}^{m-2}\binom{m}{p}\frac{(-1)^{m-p}(m-p)!}{(z-\xi_{t})^{m-p+1}}\Phi^{(p+1)}(z)+mh^{\prime}(z)\Phi_{t}^{(m)}(z)+h(z)\Phi_{t}^{(m+1)}(z).$
Consider the Taylor expansion of each derivative of $\Phi_{t}$ about $\xi_{t}=e^{i\phi_{t}^{j}}$,
$\Phi_{t}^{(p+1)}(z)$ $=\Phi_{t}^{(p+1)}(\xi_{t})+(z-\xi_{t})\Phi_{t}^{(p+2)}(\xi_{t})+\dots+\frac{(z-\xi_{t})^{m-p+1}}{(m-p+1)!}\Phi_{t}^{(m+2)}(\xi_{t})$
$\phantom{=}+O\left((z-\xi_{t})^{m-p+2}\sup_{|w-\xi_{t}|<|z-\xi_{t}|}|\Phi_{t}^{(m+3)}(w)|\right).$
Substituting each of these into (4.2), for $11\leq r\leq m-$ 1, the coefficient of $\Phi_{t}^{(r)}(\xi_{t})$ is
$2\xi_{t}^{2}\sum_{p=0}^{r-1}\binom{m}{p}\frac{(-1)^{m-p}(m-p)!}{(z-\xi_{t})^{m-p+1}}\frac{(z-\xi_{t})^{r-p-1}}{(r-p-1)!}$ $=\frac{2\xi_{t}^{2}}{(z-\xi_{t})^{m-r+2}}\sum_{p=0}^{r-1}\frac{m!}{p!(r-1-p)!}(-1)^{m-p}$
$=\frac{2\xi_{t}^{2}\,m(m-1)\dots r}{(z-\xi_{t})^{m-r+2}}\sum_{p=0}^{r-1}\binom{r-1}{p}(-1)^{m-p}.$
If $r>1$ this is just a multiple of the binomial expansion of $(1-1)^{r-1}=0$, so there is no $\Phi_{t}^{(r)}(\xi_{t})$ term in $q_{t}^{(m)}(\xi_{t})$. If $r=1$ the coefficient is non-zero, but $\Phi_{t}^{\prime}(\xi_{t})=0$. Next, the coefficient of $\Phi_{t}^{(m)}(\xi_{t})$ is, again comparing coefficients with the binomial expansion of $(1-1)^{m-1}$,
$2\xi_{t}^{2}\sum_{p=0}^{m-2}\binom{m}{p}\frac{(-1)^{m-p}(m-p)!}{(z-\xi_{t})^{m-p+1}}\frac{(z-\xi_{t})^{m-p-1}}{(m-p-1)!}+m(1-\frac{2\xi_{t}^{2}}{(z-\xi_{t})^{2}})$
$=\frac{2m\xi_{t}^{2}}{(z-\xi_{t})^{2}}\sum_{p=0}^{m-2}\binom{m-1}{p}(-1)^{m-p}+m(1-\frac{2\xi_{t}^{2}}{(z-\xi_{t})^{2}})$
$=\frac{2m\xi_{t}^{2}}{(z-\xi_{t})^{2}}(-(1-1)^{m-1}+1)+m(1-\frac{2\xi_{t}^{2}}{(z-\xi_{t})^{2}})$
$=m$
if $m\geq 2$. Then writing $\sum_{p=0}^{m-2}\binom{m}{p}(-1)^{m-p}=(1-1)^{m}-\binom{m}{m-1}(-1)^{1}-\binom{m}{m}(-1)^{0}=m-1$, we find that the coefficient of $\Phi_{t}^{(m+1)}(\xi_{t})$ in the expansion of (4.2) is
$2\xi_{t}^{2}\sum_{p=0}^{m-2}\binom{m}{p}\frac{(-1)^{m-p}(m-p)!}{(z-\xi_{t})^{m-p+1}}\frac{(z-\xi_{t})^{m-p}}{(m-p)!}+(m-\frac{2m\xi_{t}^{2}}{(z-\xi_{t})^{2}})(z-\xi_{t})+z+2\xi_{t}+\frac{2\xi_{t}^{2}}{z-\xi_{t}}$
$=\frac{2\xi_{t}^{2}}{z-\xi_{t}}(0-\binom{m}{m-1}(-1)^{1}-\binom{m}{m}(-1)^{0})+m(z-\xi_{t})-\frac{2(m-1)\xi_{t}^{2}}{z-\xi_{t}}+z+2\xi_{t}$
$=\frac{2\xi_{t}^{2}}{z-\xi_{t}}(m-1)+m(z-\xi_{t})-\frac{2(m-1)\xi_{t}^{2}}{z-\xi_{t}}+z-\xi_{t}+3\xi_{t}$
$=3\xi_{t}+(m+1)(z-\xi_{t}).$
Finally, as $\sum_{p=0}^{m-2}\binom{m+1}{p}(-1)^{m+1-p}=-\frac{1}{2}m(m+1)+m$, the coefficient of $\Phi_{t}^{(m+2)}(\xi_{t})$ is
$2\xi_{t}^{2}\sum_{p=0}^{m-2}\binom{m}{p}\frac{(-1)^{m-p}(m-p)!}{(z-\xi_{t})^{m-p+1}}\frac{(z-\xi_{t})^{m-p+1}}{(m-p+1)!}$
$\phantom{=}+\frac{m}{2}(1-\frac{2\xi_{t}^{2}}{(z-\xi_{t})^{2}})(z-\xi_{t})^{2}+(z+2\xi_{t}+\frac{2\xi_{t}^{2}}{z-\xi_{t}})(z-\xi_{t})$
$=2\xi_{t}^{2}\sum_{p=0}^{m-2}\frac{m!}{p!(m-p+1)!}(-1)^{m-p}-m\xi_{t}^{2}+2\xi_{t}^{2}+O(z-\xi_{t})$
$=\xi_{t}^{2}\left(\frac{-2}{m+1}\sum_{p=0}^{m-2}\binom{m+1}{p}(-1)^{m+1-p}-m+2\right)+O(z-\xi_{t})$
$=\xi_{t}^{2}\left(\frac{2}{m+1}(\frac{(m+1)m}{2}-m)-m+2\right)+O(z-\xi_{t})$
$=\frac{2\xi_{t}^{2}}{m+1}+O(z-\xi_{t}),$
and so taking the limit $z\to\xi_{t}$, we have
$q_{t}^{(m)}(\xi_{t})$ $=m\Phi_{t}^{(m)}(\xi_{t})+3\xi_{t}\Phi_{t}^{(m+1)}(\xi_{t})+\frac{2\xi_{t}^{2}}{m+1}\Phi_{t}^{(m+2)}(\xi_{t}),$
as claimed. ∎
Corollary 4.6.
When $\xi_{t}\not=x_{j}(t)$, we have $q_{t}(x_{j}(t))=\Phi_{t}^{\prime}(x_{j}(t))\,x_{j}(t)\frac{x_{j}(t)+\xi_{t}}{x_{j}(t)-\xi_{t}}=0$, and so for all $t$,
$q_{t}(x_{j}(t))=2I^{j}_{t}x_{j}(t)^{2}\Phi_{t}^{\prime\prime}(x_{j}(t)).$
By a similar argument, if $\bar{q}_{t}(z)=\bar{\Phi}_{t}^{\prime}(z)z\sum_{j=1}^{k}\bar{p}^{j}_{t}\frac{z+\bar{x}_{j}(t)}{z-\bar{x}_{j}(t)}$, then
$\bar{q}_{t}(\bar{x}_{j}(t))$ $=2\bar{p}^{j}_{t}\bar{x}_{j}(t)^{2}\bar{\Phi}_{t}^{\prime\prime}(\bar{x}_{j}(t)).$
Proof of Proposition 4.4.
Equation (2.7) of [1] gives a useful expression for the evolution of $e^{i\phi^{j}_{t}}$ over time:
@@
and so as the denominator is bounded below by $L=L(k,K_{0},T,\eta)$,

### v3

@@ 4.1 ALE to auxiliary model
The two maps above whose derivatives we consider can be generated by the backward equation using driving functions which we will call $\xi^{0}$ (generating $\Phi^{\mathrm{ALE}}_{0}\circ\dots\circ\Psi_{N_{j}-1}^{j}$) and $\xi^{1}$ (generating $R_{\Delta_{1}^{n}}\circ\dots\circ\overline{\Psi}^{j}_{N_{j}-1}$). We can now use Lemma 3.5 to compare the above derivatives. We have the bound $\|\xi^{0}-\xi^{1}\|_{\infty}\leq|\delta_{1}|+\dots+|\delta_{n}|$. Taking $z_{0}=f^{\mathrm{ALE}}_{n_{j}(N_{j})}(\Psi_{N_{j}}(e^{\sigma+i\theta}e^{i\phi_{n}^{j}}))$, we have $|z_{0}|-1\geq\mathbf{c}^{1/2}$, so using standard estimates for conformal maps [3, Chapter 4.4, Corollary 6] we have $\frac{\mathbf{c}^{1/2}}{A}\leq|(u_{t}^{0})^{\prime}(z_{0})|\leq\frac{A}{\mathbf{c}^{1/2}}$ for all $0<t<T$, where the constant $A$ depends only on $T$ and $K_{0}$. This gives us, modifying the constant $A=A(T,K_{0})$ appropriately (as we will throughout), $\frac{\mathbf{c}^{1/2}}{A}t\leq|\Lambda_{t}|\leq\frac{A}{\mathbf{c}^{3/2}}t$.
@@
$f(w)=1+d(\mathbf{c})+O\left(\frac{|w-1|^{2}}{\mathbf{c}^{1/2}}+|w-1|\right),$ (4.8)
@@
(see [1, equation (26)]) and using (4.5) and (4.8) we know
@@
and so by standard conformal map estimates (see [3, Chapter 4.4]),
@@
Thus, using the bound $|e^{\sigma+i\theta}-1|^{2}=(e^{\sigma}-1)^{2}+2e^{\sigma}(1-\cos\theta)\leq A(\sigma^{2}+\theta^{2})$ for $|\theta|<\mathbf{c}$ and substituting $x=\theta/\sigma$, we get
@@
[…]
Case I: First suppose there is no such $k$. Then by Lemma 4.6, $|(f^{\mathrm{ALE}}_{1}\circ\dots\circ f^{\mathrm{ALE}}_{n})^{\prime}(\hat{z})|\leq B\mathbf{c}^{-1/2}$ almost surely, so by (3.7) from Lemma 3.5,
@@
[…]
@@
Case II: If there is only one such $k$, we must have $|\hat{z}_{k}-e^{i\theta^{\mathrm{ALE}}_{k-1}}|\geq\frac{\mathbf{c}^{1/2}}{2A_{T}^{2}},$ since there is an $l\geq k$ with $\theta^{\mathrm{ALE}}_{l}$ attached at the same slit as $\theta^{\mathrm{ALE}}_{k-1}$ and if $\hat{z}_{k}$ were any closer to $e^{i\theta^{\mathrm{ALE}}_{k-1}}$ then $\hat{z}_{l+1}$ would be within $|e^{i(\beta_{\mathbf{c}}+\frac{\mathbf{c}^{1/2}}{A_{T}})}-1|$ of $e^{i\theta^{\mathrm{ALE}}_{l}}$.
@@
and Lemma 3.3 tells us that $|z_{k-1}-e^{i\theta^{\mathrm{ALE}}_{k-1}}|\asymp\mathbf{c}^{1/4}\delta^{1/2}$. If the attachment point $\theta^{\mathrm{ALE}}_{k-1}$ is on the $j$th slit, then it is within $D$ of the preimage $\phi^{j}_{k-2}$ of $\theta^{\mathrm{ALE}}_{l}$ under $f^{\mathrm{ALE}}_{l+1}\circ\dots\circ f^{\mathrm{ALE}}_{k-2}$, where l<k1l<k-1 was the previous time a particle was attached at the $j$th slit (without loss of generality such an $l$ exists, as we can decompose the initial condition to have a slit of capacity $\mathbf{c}$ at the top of a longer slit in position $j$).
@@
Case III: If there are two values of $k$, $k_{2}<k_{1}$, we can find a bound of the same order using the same argument as above, replacing $k$ by $k_{1}$ and $l$ by $k_{2}$. ∎
@@ 4.2 Auxiliary to multinomial model
$\left|\log\frac{|(\Phi^{*}_{l})^{\prime}(e^{\sigma+i(\phi_{l}^{j}+\theta)})|}{|e^{\sigma+i\theta}-1|\big|(\Phi^{*}_{l})^{\prime\prime}(e^{i\phi_{l}^{j}})\big|}\right|$ $\leq\frac{|e^{\sigma+i\theta}-1|\sup_{z\in\gamma}|(\Phi^{*}_{l})^{(3)}(z)|}{|(\Phi^{*}_{l})^{\prime\prime}(e^{i\phi_{l}^{j}})|}.$ (4.18)
@@ 5 Multinomial to Laplacian path model
where $\mathcal{H}=\{\varphi\in C(X):\|\varphi\|_{\mathrm{Lip}}+\|\varphi\|_{\infty}\leq 1\}$, for
@@
Under the same assumptions as in Theorem 2.1, there exists a universal constant $R$ and a constant $A=A(T,L,k,K_{0},\eta)$ such that
@@
To derive Corollary 5.3 from Theorem 5.2, we also use Proposition 5.4, which we will prove in the following section.
@@
The latter two terms are small with high probability using Theorem 5.2, so we only need to bound the “martingale” terms $\int_{0}^{T}(I^{j}_{t}-p^{j}_{t})\varphi(x_{j}(t))\,\mathrm{d}t$ for each $j$. We can write each term as
@@
Since $|p^{j}_{n\mathbf{c}}-p^{j}_{t}|\leq A\mathbf{c}$ by Proposition 5.4, we have $|\sum_{n=0}^{\frac{T}{\mathbf{c}}-1}\int_{n\mathbf{c}}^{(n+1)\mathbf{c}}(p^{j}_{n\mathbf{c}}-p^{j}_{t})\varphi(x_{j}(t))\,\mathrm{d}t|\leq AT\mathbf{c}$. For the remaining term, we would like to simply compute second moments and so show the term is small, but need some way of doing this uniformly in $\varphi$. Choose a large $N$, and for simplicity assume that $T/\mathbf{c}$ is an integer multiple of $N$. Then write
@@
Our proof of Theorem 5.2 has two parts: first we show in Proposition 5.4 that $t\mapsto(x_{j}(t),\bar{x}_{j}(t),p^{j}_{t},\bar{p}^{j}_{t})$ is Lipschitz on $[0,T]$, which will allow us to deduce global bounds on $|x_{j}(t)-\bar{x}_{j}(t)|$ and $|p^{j}_{t}-\bar{p}^{j}_{t}|$ by showing that they are close at a sufficiently dense set of times.
@@ 5.1 Lipschitz bound on the locations and probabilities
There exists a constant $A=A(k,K_{0},T,\eta)$ such that, almost surely, for $0\leq t,s\leq T$,
$|x_{j}(t)-x_{j}(s)|\leq A|t-s|\quad\text{ and }\quad|p_{t}^{j}-p_{s}^{j}|\leq A|t-s|,$
for all $j=1,\dots,k$ in the multinomial model. For the Laplacian path model, for all $j$ and for $0\leq t,s\leq T$, we also have
$|\bar{x}_{j}(t)-\bar{x}_{j}(s)|\leq A|t-s|\quad\text{and}\quad|\bar{p}_{t}^{j}-\bar{p}_{s}^{j}|\leq A|t-s|.$
@@
Equation (2.7) of [2] gives a useful expression for the evolution of $x_{j}(t)$ over time:
@@
and so as the denominator is bounded below by $L=L(k,K_{0},T,\eta)$, we have